OpenStudy (anonymous):

12. Find the least common multiple of y2 – 2y – 15 and y2 – 7y + 10. (1 point) (y – 2)(y + 3)(y – 5) (y + 2)(y + 3)(y – 5) (y – 2)(y – 3)(y – 5) (y + 2)(y + 3)(y + 5)

OpenStudy (mathstudent55):

First, factor each polynomial.

OpenStudy (whpalmer4):

To find the LCM of two things, you need to factor each one. When you have factored them, the LCM is the product of the highest power of each factor found. For example, the LCM of 25 and 45: 25 = 5*5 45 = 3*3*5 The LCM of 25 and 45 is 3*3*5*5 = 9*25 = 225 To check, write out a list of multiples of each number: 25 50 75 100 125 150 175 200 225 45 90 135 180 225 225 is the first number that appears in both lists, so it is the least common multiple. You do the same thing with algebraic expressions. Find the LCM of x^2 - 2x + 1 and x^2 - 3x + 2: x^2 - 2x + 1 factors to (x-1)(x-1) x^2-3x+2 factors to (x-1)(x-2) The LCM is therefore (x-1)(x-1)(x-2)

OpenStudy (e.cociuba):

Ans is a) (y - 2)(y + 3)(y - 5)

OpenStudy (anonymous):

How did you get that e?

OpenStudy (e.cociuba):

y^2 – 2y – 15 = y^2 + 3y - 5y - 15 = y(y + 3) - 5(y + 3) = (y + 3)(y - 5) y^2 - 7y + 10 = y^2 - 5y - 2y + 10 = y(y - 5) - 2(y - 5) = (y - 5)(y - 2) Least common factor = (y + 3)(y - 2)(y - 5) Therefore . Ans is a) (y - 2)(y + 3)(y - 5)

OpenStudy (anonymous):

Thank you guys! Especcially e and Whpalmer

OpenStudy (e.cociuba):

np:)